OpenStudy (adellicolli01):

Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). (1 point)

OpenStudy (1davey29):

First, find g(2), g(2)=2^2-6(2)-7=-15. Then plug in g(2) into your f(x), f(g(2)=f(-15)=-15+8=-7.

OpenStudy (mhchen):

Plug in g(x) inside f(x) so g(2) = 2^2 - 6(2) - 7 = something Then plug that something into f(x) and you get something + 8.